The directions are:
Factor each polynomial, if possible. If the polynomial cannot be factored using intergers, write prime.
The problem is:
12q^2+34q-28

Answer by edjones(8007)

12q^2+34q-28
=2(6q^2+17q-14)
6*-14=-84
2 factors of -84 whose sum is 17 are 21 and -4
2(6q^2+21q-4q-14)
=2(3q(2q+7)-2(2q+7))
=2(3q-2)(2q+7)
.
Ed